# M2D2: Exploring General-purpose Audio-Language Representations Beyond CLAP

### v1

@@ IV-C2 Audio-Text Retrieval
We evaluated the effectiveness of the CLAP features in the audio-text retrieval task setting. This task involves searching for text from audio or audio from text, so unlike zero-shot classification, it requires fine-grained comparisons to find pairs at the data instance level[80]. Therefore, the more accurately the features represent the audio and text, the better the task performance.
@@
Table VIII shows that M2D2- improves overall performance compared to M2D20- and M2D20-+WC, which is pre-trained on the same dataset as M2D2- to make the difference only in the text encoder and audio projector used. M2D2- improves performance on most metrics, e.g., AudioCaps Text-to-Audio R@1 performance of 27.34 compared to M2D20-+WC’s 26.06. The results confirm the effectiveness of the M2D2- updates from M2D20- (M2D-CLAP) using an LLM-based text (sentence) encoder and transformer-based audio projector.
M2D2 further improves overall performance compared to M2D2-, by pre-training the text encoder in the second stage. For example, the AudioCaps R@1 performance improved from M2D2-’s 27.34 to M2D2’s 37.24. The results demonstrate the importance of training the text encoder to achieve fine-grained alignment of the paired features.
The results in Table VIII indicate that the combination of datasets significantly impacts performance. The “Final CLAP training dataset” column in the table lists the datasets used in the model’s last training just before the evaluation. LAION-CLAP, CLAP2023, and M2D2, which used a large-scale dataset (denoted as ”LS” in the table) just before the evaluation, perform poorly compared to those without one. Their AudioCaps Text-to-Audio R@1 performance is about 35 to 37, which is inferior to the 39.7 of HTSAT-BERT-PT without using large-scale datasets, for example. These results indicate that a limited combination of WavCaps, AudioCaps, and Clotho is advantageous for evaluation in the benchmark. As an exception, Cacophony, which developed its data curation strategy, performs well with large-scale datasets.
M2D2+ and M2D2 achieve top performance when additionally trained under conditions advantageous to the benchmark as in the previous studies[4][6][7]. M2D2+ outperforms or becomes comparable to HTSAT-BERT-PT after pre-training on the same data (WavCaps, AudioCaps, and Clotho) as HTSAT-BERT-PT in additional stage 2.1. M2D2 even outperforms top performers HTSAT-BERT-FT, Cacophony, and FLAP in AudioCaps after fine-tuning its audio encoder to AudioSet in the additional stage 1.1.
In summary, the experiments validated the effectiveness of M2D2. M2D2 performs comparably to previous studies, and it further improves performance to the top level through additional stages for a fair comparison to previous studies.
@@
| (Proposed models) | |||||||||||
| M2D2- | - | 23.3 0.3 | 68.3 1.5 | 16.9 0.2 | 42.6 0.8 | 61.9 0.5 | 17.4 0.1 | 41.2 0.3 | 12.2 0.1 | 26.7 0.2 | 48.9 0.2 |
| M2D2 | - | 23.7 0.3 | 69.6 1.3 | 17.2 0.3 | 43.4 0.7 | 62.5 0.4 | 17.8 0.1 | 43.5 0.4 | 12.4 0.1 | 28.0 0.2 | 49.9 0.2 |
| M2D2+ | - | 23.6 0.3 | 69.4 1.4 | 17.1 0.3 | 43.2 0.8 | 62.3 0.5 | 17.8 0.1 | 43.2 0.4 | 12.5 0.1 | 27.8 0.3 | 49.7 0.2 |
| M2D2 | ✓ | 24.3 0.4 | 72.4 1.8 | 17.6 0.3 | 45.0 1.0 | 63.9 0.4 | 17.7 0.1 | 43.1 0.7 | 12.5 0.2 | 27.8 0.4 | 49.6 0.4 |
@@ IV-C3 Audio Captioning
We conducted evaluations on AudioCaps and Clotho and used a learning rate of for AudioCaps and for Clotho. We used the same settings for other hyperparameters as in EnCLAP, such as training epochs of 15. We tested the top-five checkpoints of SPIDEr validation performance on the test set and averaged the results to obtain the final result. The evaluation metrics are METEOR[83], CIDEr[84], SPICE[85], SPIDEr[86], and FENSE[87].
@@
Experimental results in Table IX show that M2D2 is comparable to LAION-CLAP on AudioCaps, while M2D2, M2D2+, and M2D2 outperform LAION-CLAP on Clotho. We also evaluated CLAP2023; however, it underperformed LAION-CLAP. Note that the experiments used the same EnCLAP codebase; the only difference is the CLAP audio feature used by the EnCLAP.
The AudioCaps results demonstrate that fine-tuning the audio encoder to AudioSet is vital for good performance; M2D2 performs comparably to LAION-CLAP while M2D2+ does not. On the other hand, the Clotho results suggest that fine-tuning to AudioSet does not offer general-purpose performance; M2D2 and M2D2+ perform comparably to M2D2, and they all outperform LAION-CLAP. Notably, M2D2 performs slightly better than M2D2+ and M2D2 on Clotho, demonstrating that it learns generalizable CLAP features without the need for benchmark techniques.
In summary, the experimental results confirm that the M2D2 models achieve CLAP features that also perform well in audio captioning.
IV-D M2D2 Analysis
@@ IV-D1 Taming the CLAP objective in the first stage
When the weight is at 0.1, (b) ESC-50/US8K performance matures quickly and degrades as training progresses, while (c) performance for other tasks does not improve. When the weight is at 0.001, (b) ESC-50/US8K performance improves slowly, suggesting the weight is too low. In contrast, with the weight at 0.01, performance for all tasks matures as training progresses.
The ESC-50/US8K performance’s maturing quickly with high weight is similar to the observation that the CLAP training matures early in the previous study[6]. The observations of all cases suggest that combining SSL and the CLAP objectives while suppressing CLAP is essential for successful pre-training.
@@
The audio projector, which aggregates the audio features of each token into a single feature, impacts the effectiveness of training through CLAP loss calculation. Fig. 5 shows the average results of linear evaluation across training epochs for the ablations of audio projectors: M2D2’s one-block transformer encoder, the two-block encoder (one additional block), and one-hidden-layer MLP used in M2D-CLAP (M2D20-). When using MLP as a projector, we averaged all the audio encoder output features and fed the averaged feature to it.
The results in Fig. 5(a) show that MLP exhibits the best results up to 200 epochs but stalls afterward. In the final 300 epochs, one block transformer of M2D2 outperforms the others and obtains the best results. The MLP results in Fig. 5(b) show a trend similar to the ESC-50/US8K results when the CLAP loss weight is high () in Fig. 4(b), suggesting that the use of a simple MLP augments the contribution of the CLAP objective during training and leads to early maturation, as reported in the previous study[6].
The two-block transformer underperformed the one-block transformer throughout the training process, indicating that using a projector with adequate network capacity is also vital for successful pre-training.
@@ IV-D2 Text encoder ablations
We assessed the performance impact on the type of text encoder used in the first stage. We compared the LLM-based model NV-Embed-v2[21][42] (7.85B parameters) used in M2D2 against the GTE-Base[60] (109M parameters) used in M2D-CLAP (M2D20-) and GTE-Qwen2-7B-instruct999https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct (7.61B parameters)[60], which ranked higher than NV-Embed-v2 on the MTLB leaderboard. The experiments replaced the text encoder in M2D2. We also assessed the performance of the second-stage text encoder when using GTE-Base instead of BERT.
The linear evaluation results on the general audio tasks in Table X and the audio-text retrieval of the CLAP features in Table XI show that using NV-Embed-v2’s LLM-based sentence features achieves the best performance with both features. With a non-LLM-based GTE-Base, M2D2 learns competitive audio features while underperforming second-stage CLAP features compared to LLM-based encoders. The GTE-Qwen2-7B-instruct’s results are comparable to (while slightly worse than) NV-Embed-v2’s, confirming the effectiveness of employing LLM-based text encoders.
@@
Notably, the performance boost in the second-stage CLAP features indicates that the semantic knowledge accumulated in the audio features effectively guides M2D2 to learn useful CLAP features. The second-stage results of LLM-based NV-Embed-v2 and GTE-Qwen2-7B-instruct are 37.24 and 36.78 for R@1 of AudioCaps T-to-A, while the non-LLM-based GTE-Base’s result is 34.04, for example.
@@ IV-D3 Training design ablations
Training audio encoder in the second stage
While M2D2 freezes the audio encoder in the second stage, an alternative would be to pre-train the audio and text encoder together in this stage, as in [10]. We performed an ablation experiment in which we pre-trained the audio encoder in the second stage. In this case, computational resource constraints decreased the batch size from 2048 to 1024. We also compared the results with M2D2, in which we fine-tuned the audio encoder to AudioSet after the first stage, i.e., the additional stage 1.1 described in Section III-B4.
The results for the audio features shown in Table XII indicate that M2D2, which trains the audio encoder in the second stage, improves the performance of the audio features in the ESC-50 (ESC) and US8K tasks while degrading the performance in other tasks. This trend is the same as M2D2.
On the other hand, Table XIII shows that the audio-text retrieval performance of M2D2 is comparable to M2D2, indicating that training the audio encoder in the second stage does not improve the CLAP feature, which reveals the difference from M2D2. In addition, the decreased batch size may have affected the results.
These experiments suggest that training audio encoders in the second stage, as in the previous study Cacophony[10], might not be a good pre-training strategy for learning representations that perform well for both audio and CLAP features, the goal of M2D2.
@@
| M2D2, M2D2- | 95.2 | 89.5 | 95.6 | 68.6 | 98.1 | 73.4 | 86.6 | 76.5 | 42.2 | 80.6 |
| M2D2, M2D2 | 97.9 | 89.7 | 94.8 | 56.3 | 97.1 | 69.7 | 86.3 | 76.7 | 41.1 | 78.8 |
| (Section IV-D3: Jointly training text/audio encoders in the second stage) | ||||||||||
| M2D2 | 97.0 | 90.1 | 93.3 | 60.4 | 96.7 | 71.1 | 86.3 | 76.1 | 42.0 | 79.0 |
| (Section IV-D3: Text projector only) | ||||||||||
| M2D2-(TP only) | 94.9 | 88.0 | 94.0 | 60.5 | 97.5 | 71.6 | 83.7 | 76.0 | 41.9 | 78.7 |
| (Section IV-D3: Both audio and text projector, as in previous CLAP models) | ||||||||||
| M2D2-(+TP) | 96.1 | 88.9 | 95.6 | 66.2 | 97.8 | 72.7 | 84.7 | 77.6 | 42.3 | 80.2 |
@@
| M2D2 | 37.24 | 52.00 | 52.56 | 38.74 | 18.09 | 28.51 | 24.88 | 15.53 |
| M2D2 | 39.06 | 54.29 | 54.02 | 39.95 | 18.30 | 29.15 | 22.78 | 15.62 |
| (Section IV-D3: Jointly training text/audio encoders in the second stage) | ||||||||
| M2D2 | 37.81 | 52.29 | 50.78 | 37.15 | 17.76 | 28.83 | 21.44 | 14.11 |
| (Section IV-D3: Text projector only) | ||||||||
| M2D2 (TP only) | 34.04 | 48.73 | 45.98 | 34.42 | 17.07 | 27.44 | 22.11 | 14.56 |
| (Section IV-D3: Both audio and text projector, as in previous CLAP models) | ||||||||
| M2D2 (+TP) | 36.07 | 50.52 | 48.69 | 36.93 | 17.53 | 28.09 | 22.58 | 14.72 |
Using text projector only
In this ablation experiment, we removed the audio projector and attached a text projector on top of the text encoder to test the case where an audio feature is learned in the semantic space, and we mapped text features to the space, i.e., an audio feature was learned that works as both conventional and CLAP audio features. We used a one-hidden layer MLP as the text projector.
Adding text projector
This ablation experiment added a text projector to the text encoder, making M2D2 follow a typical CLAP configuration with both audio and text projectors. The results in Tables XII and XIII show that M2D2 (+TP), text projector-added model, degrades performance from M2D2 as in the text projector only ablation. Considering these results, we do not employ a text projector in M2D2.
@@ IV-D4 Visualizations
We visualize the pre-trained CLAP audio features and show their impact on two downstream tasks with an inherently different nature. We encoded ESC-50 and VoxCeleb1 sounds using M2D2, M2D2 (M2D2 fine-tuned to AudioSet), and CLAP and visualized their features using t-SNE as shown in Fig. 6. We use all 50 class samples from ESC-50 and samples of random ten speakers from VoxCeleb1.
The ESC-50 results show that for all audio features, the samples are clustered by class, indicating that the features are useful for the 50-class classification task. On the other hand, the VoxCeleb1 results show that the CLAP clusters have unclear boundaries compared to M2D2 and M2D2. These observations support the VC1 performance difference in the linear evaluation shown in Table II in Section IV-B1.
The CLAP features are advantageous for ESC-50 since it is an environmental sound classification task, and the pre-training captions contain much information indicating differences in environmental sounds. On the other hand, VoxCeleb1 is a 1251-speaker identification task, and the pre-training captions cannot contain speaker names or IDs, thus making it challenging for conventional CLAP features in the identification task. The conventional CLAP model would learn to represent speaker attributes such as adult or child described in captions, which may explain the tendency not to cluster by individuals. In contrast, M2D2 features form more distinct clusters, indicating that SSL helps learn more information about sound itself, which leads to clearer boundaries for individuals.
These observations demonstrate that jointly learning SSL and CLAP allows M2D2 to include diverse information in CLAP audio features.
@@
Figure 7 shows the attention in the audio projector for the audio samples from AudioCaps and Clotho. The attention map directly shows which patch features the final CLAP audio feature attends to because the audio projector is a single-layer transformer encoder with a single head. We used the M2D2 model and picked the sounds with typical attention maps.

### v2

@@ IV-C2 Audio-Text Retrieval
The audio-text retrieval task evaluates CLAP features and involves searching for text from audio or audio from text, so unlike zero-shot classification, it requires fine-grained comparisons to find pairs at the data instance level[82]. Therefore, task performance improves as the features more accurately align the audio and text.
@@
Table VIII shows that M2D-CLAP achieves top performance, and the ablation variants improve the performance as stages go by. The first-stage model, M2D-CLAP, exhibits a large gap from the previous SOTA results, such as HTSAT-BERT-FT. The second-stage pre-training significantly improves the performance, e.g., from 27.34 of the first stage to 37.24 of the second stage for AudioCaps Text-to-Audio R@1. This indicates that pre-training the text encoder is crucial for a task that requires finer alignment between audio and text, unlike the zero-shot classification, where the first-stage model achieved comparable results. The additional stages 1.1 (AudioSet fine-tuning) and 2.1 (WavCaps fine-tuning) bring the final performance to the top level, and M2D-CLAP even outperforms top performers HTSAT-BERT-FT, Cacophony, and FLAP in AudioCaps.
M2D-CLAP outperforms both M2D-CLAP and M2D-CLAP+WC. Notably, the latter is pre-trained on the same dataset as M2D-CLAP, ensuring that the only differences lie in the text encoder and audio projector. M2D-CLAP improves performance on most metrics, e.g., AudioCaps Audio-to-Text R@1 performance of 36.05 compared to M2D-CLAP+WC’s 28.63. The results confirm the effectiveness of using an LLM-based text (sentence) encoder and a transformer-based audio projector.
Table VIII indicates that the combination of datasets significantly impacts performance. The “Final CLAP training dataset” column lists the datasets used in the model’s last training just before the evaluation. LAION-CLAP, CLAP2023, and M2D-CLAP, which used a large-scale dataset (denoted as ”LS” in the table), perform poorly compared to those without one. Their AudioCaps Text-to-Audio R@1 performance is about 35 to 37, which is inferior to the 39.7 of HTSAT-BERT-PT without using large-scale datasets, for example. These results indicate that a limited combination of WavCaps, AudioCaps, and Clotho is advantageous for evaluation in the benchmark. As an exception, Cacophony, which developed its data curation strategy, performs well with large-scale datasets.
In summary, the experiments validated the effectiveness of M2D-CLAP and demonstrated that the staged training process enables the model to achieve fine-grained alignment between paired features.
@@
| (Proposed model) | |||||||||||
| M2D-CLAP | ✓ | 24.3 0.4 | 72.4 1.8 | 17.6 0.3 | 45.0 1.0 | 63.9 0.4 | 17.7 0.1 | 43.1 0.7 | 12.5 0.2 | 27.8 0.4 | 49.6 0.4 |
| M2D-CLAP | - | 23.3 0.3 | 68.3 1.5 | 16.9 0.2 | 42.6 0.8 | 61.9 0.5 | 17.4 0.1 | 41.2 0.3 | 12.2 0.1 | 26.7 0.2 | 48.9 0.2 |
| M2D-CLAP | - | 23.7 0.3 | 69.6 1.3 | 17.2 0.3 | 43.4 0.7 | 62.5 0.4 | 17.8 0.1 | 43.5 0.4 | 12.4 0.1 | 28.0 0.2 | 49.9 0.2 |
| M2D-CLAP | - | 23.6 0.3 | 69.4 1.4 | 17.1 0.3 | 43.2 0.8 | 62.3 0.5 | 17.8 0.1 | 43.2 0.4 | 12.5 0.1 | 27.8 0.3 | 49.7 0.2 |
@@ IV-C3 Audio Captioning
We conducted evaluations on AudioCaps and Clotho and used a learning rate of for AudioCaps and for Clotho. We used the same settings for other hyperparameters as in EnCLAP, such as training epochs of 15. We tested the top five checkpoints of SPIDEr validation performance on the test set and averaged the results to obtain the final result. The evaluation metrics are METEOR[85], CIDEr[86], SPICE[87], SPIDEr[88], and FENSE[89].
@@
Table IX shows that M2D-CLAP archives top performance, which is comparable to LAION-CLAP on AudioCaps and better than LAION-CLAP on Clotho. We also evaluated CLAP2023; however, it underperformed LAION-CLAP. Note that the experiments used the same EnCLAP codebase; the only difference is the CLAP audio feature used by EnCLAP.
The AudioCaps results of intermediate stages, M2D-CLAP, underperform those of LAION-CLAP, indicating that fine-tuning the audio encoder to AudioSet is vital for good performance with AudioCaps. On the other hand, the Clotho results suggest that fine-tuning to AudioSet does not offer general-purpose performance; all M2D-CLAP variants outperform LAION-CLAP.
Notably, M2D-CLAP performs slightly better than M2D-CLAP on Clotho, demonstrating that it learns generalizable CLAP features without the need for fine-tuning to AudioSet and WavCaps.
In summary, the experimental results confirm that M2D-CLAP produces CLAP features that also perform well in audio captioning.
IV-D M2D-CLAP Analysis
@@ IV-D1 Taming the CLAP objective in the first stage
With a weight of 0.1, (b) ESC-50/US8K performance matures quickly but degrades later, while (c) other tasks show no improvement. At 0.001, (b) performance improves slowly, indicating the weight is too low. In contrast, 0.01 allows (a) all tasks to mature steadily during training. For reference, at weight 0.0 (i.e., M2D), (b) is low due to lack of caption supervision, (c) remains high, and (a) is slightly lower overall than M2D-CLAP.
The rapid maturation of ESC-50/US8K performance with higher CLAP weights mirrors the early maturation of CLAP training reported in prior work[6]. These results suggest that suppressing the effect of CLAP while combining it with SSL is key for effective pre-training.
@@
The audio projector, which aggregates token-level audio features into a single feature for CLAP loss computation, impacts training effectiveness. Fig. 5 compares average linear evaluation results across epochs for different audio projectors: the one-block transformer in M2D-CLAP, a two-block transformer (with one additional block), and a one-hidden-layer MLP used in the previous M2D-CLAP[11]. For the MLP, we input the average of all audio encoder outputs.
In Fig. 5(a), the MLP performs best up to 200 epochs but stalls afterward. In the final 300 epochs, M2D-CLAP’s one-block transformer surpasses the others. The MLP’s trend (Fig. 5(b)) resembles ESC-50/US8K results under a high CLAP loss weight () in Fig. 4(b), suggesting that a simple MLP increases CLAP’s influence during training and leads to early maturation, in line with previous findings[6].
The two-block transformer consistently underperforms the one-block version, indicating that moderate network capacity is important for effective pre-training.
@@ IV-D2 Text encoder ablations
We assessed the performance impact of the type of text encoder used in the first stage. We compared the LLM-based model NV-Embed-v2[21][42] (7.85B parameters) used in M2D-CLAP against the GTE-Base[64] (109M parameters) used in the previous M2D-CLAP (M2D-CLAP) and GTE-Qwen2-7B-instruct999https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct (7.61B parameters)[64], which ranked higher than NV-Embed-v2 on the MTLB leaderboard. The experiments replaced the text encoder in M2D-CLAP. We also assessed the performance of the second-stage text encoder when using GTE-Base instead of BERT.
The linear evaluation results on the general audio tasks in Table X and the audio-text retrieval of the CLAP features in Table XI show that using NV-Embed-v2’s LLM-based sentence features achieves the best performance with both features. With a non-LLM-based GTE-Base, M2D-CLAP learns competitive audio features while underperforming second-stage CLAP features compared to LLM-based encoders. The GTE-Qwen2-7B-instruct’s results are comparable to (while slightly worse than) NV-Embed-v2’s, confirming the effectiveness of employing LLM-based text encoders.
@@
Notably, the performance boost in the second-stage CLAP features indicates that the semantic knowledge accumulated in the audio features effectively contributes to learn useful CLAP features. The second-stage results of LLM-based NV-Embed-v2 and GTE-Qwen2-7B-instruct are 37.24 and 36.78 for R@1 of AudioCaps T-to-A, while the non-LLM-based GTE-Base’s result is 34.04, for example.
@@ IV-D3 Training design ablations
We conducted the following ablations focusing on differences from conventional CLAP in the M2D-CLAP training process:
1.
Training audio encoder in the second stage: While M2D-CLAP freezes the audio encoder in the second stage, an alternative is to pre-train both audio and text encoders, as in [10]. This ablation adopts that approach, while reducing the batch size to 1024 from 2048 due to limited computational resources.
2.
Using text projector only: This ablation removed the audio projector and added a text projector on top of the text encoder, aiming to learn an audio feature directly in the semantic space that serves both conventional and CLAP audio features. We used a one-hidden-layer MLP as the text projector.
3.
Adding text projector: This ablation added a text projector on top of the text encoder, making M2D-CLAP follow a typical CLAP configuration with both audio and text projectors.
Tables XII and XIII present the evaluation results for audio and CLAP features, respectively, with key observations summarized below:
(i) Training audio encoder in the second stage: For audio features, the ablation (Stage 2/TA) improves performance on ESC-50 (ESC) and US8K, but clearly degrades it on speech tasks. This trend, similar to that of the final M2D-CLAP, suggests an influence of caption supervision.
For CLAP features, Stage 2/TA lowers A-to-T performance in both retrieval tasks. Unlike previous studies, factors such as pre-training the audio encoder under M2D and CLAP objectives, omitting the text projector, or using a non-MLP audio projector may have slightly hindered fine-grained audio-to-text alignment. Additionally, the reduced batch size may have negatively impacted results.
(ii) Using text projector only: Unlike other ablations, the ablations (Stages 1, 2/TP only) degrade performance across all results for both features. Thus, these results highlight the challenge of bridging conventional and CLAP audio features.
(iii) Adding text projector: The ablation (Stages 1/+TP) causes minimal changes in audio feature performance, while Stages 2/+TP clearly degrades CLAP feature performance. Mapping text features through a projector appears to hinder overall audio-text alignment in CLAP features.
In summary, unlike previous work such as Cacophony[10], continuing to train the audio encoder in Stage 2 is ineffective, and incorporating a text projector yields no clear advantage for M2D-CLAP.
@@
| Stage 1 | 95.2 | 89.5 | 95.6 | 68.6 | 98.1 | 73.4 | 86.6 | 76.5 | 42.2 | 80.6 |
| (i) Training audio encoder in the second stage. | ||||||||||
| Stage 2/TA | 97.0 | 90.1 | 93.3 | 60.4 | 96.7 | 71.1 | 86.3 | 76.1 | 42.0 | 79.0 |
| (ii) Using text projector only. | ||||||||||
| Stage 1/TP only | 94.9 | 88.0 | 94.0 | 60.5 | 97.5 | 71.6 | 83.7 | 76.0 | 41.9 | 78.7 |
| (iii) Adding text projector, as in previous CLAP models. | ||||||||||
| Stage1/+TP | 96.1 | 88.9 | 95.6 | 66.2 | 97.8 | 72.7 | 84.7 | 77.6 | 42.3 | 80.2 |
| M2D-CLAP | 97.9 | 89.7 | 94.8 | 56.3 | 97.1 | 69.7 | 86.3 | 76.7 | 41.1 | 78.8 |
@@
| Stage 2 | 37.24 | 52.00 | 52.56 | 38.74 | 18.09 | 28.51 | 24.88 | 15.53 |
| (i) Training audio encoder in the second stage. | ||||||||
| Stage 2/TA | 37.81 | 52.29 | 50.78 | 37.15 | 17.76 | 28.83 | 21.44 | 14.11 |
| (ii) Using text projector only. | ||||||||
| Stage 2/TP only | 34.04 | 48.73 | 45.98 | 34.42 | 17.07 | 27.44 | 22.11 | 14.56 |
| (iii) Adding text projector, as in previous CLAP models. | ||||||||
| Stage 2/+TP | 36.07 | 50.52 | 48.69 | 36.93 | 17.53 | 28.09 | 22.58 | 14.72 |
| M2D-CLAP | 39.06 | 54.29 | 54.02 | 39.95 | 18.30 | 29.15 | 22.78 | 15.62 |
@@ IV-D4 Visualizations
We visualize the pre-trained CLAP audio features and show their impact on two downstream tasks with an inherently different nature. We encoded ESC-50 and VoxCeleb1 sounds using M2D-CLAP and CLAP and visualized their features using t-SNE as shown in Fig. 6. We use all 50 class samples from ESC-50 and samples of random ten speakers from VoxCeleb1. We also computed the silhouette score to quantitatively assess the quality of clusters formed by each model’s features.
The ESC-50 results show that both model features form clear class-wise clusters with high silhouette scores, consistent with their strong performance in the linear evaluation results in Table II in Section IV-B1. On the other hand, the VoxCeleb1 results show that CLAP forms slightly less distinct clusters (silhouette score of 0.072) compared to M2D-CLAP (silhouette score of 0.109). These observations support the VC1 performance difference in Table II.
The CLAP features are advantageous for ESC-50 since it is an environmental sound classification task, and the pre-training captions contain much information indicating differences in environmental sounds. On the other hand, VoxCeleb1 is a 1251-speaker identification task, and the pre-training captions cannot contain speaker names or IDs, thus making it challenging for conventional CLAP features in the identification task. The conventional CLAP model would learn to represent speaker attributes such as adult or child described in captions, which may explain the tendency not to cluster by individuals. In contrast, M2D-CLAP features form more distinct clusters, indicating that SSL helps learn more descriptive information about sound itself, which leads to clearer clusters for individuals.
These observations demonstrate that jointly learning SSL and CLAP allows M2D-CLAP to encode both discriminative and descriptive information in CLAP audio features.
@@
Figure 7 shows the attention in the audio projector in M2D-CLAP for the audio samples from AudioCaps and Clotho. The attention map directly shows which patch features the final CLAP audio feature attends to because the audio projector is a single-layer transformer encoder with a single head. We picked the sounds with typical attention maps.
@@
IV-D5 Potential and limitations
The application of top-performing audio features to medical (respiratory and heart) sounds in [90] demonstrates their practical utility, suggesting broader potential for real-world problem solving. CLAP features have also been applied to diverse tasks, such as music generation[91] and foley sound synthesis[92]. Thus, M2D-CLAP has the potential to contribute to a wide range of applications by providing both general audio and CLAP features.
On the other hand, it still has limitations in extending its applicability, particularly regarding time resolution and frequency range. It processes sounds with a temporal resolution of 160 ms and a sampling rate of 16 kHz, and adapting it to tasks with different requirements would necessitate retraining. While effectiveness has been shown even with limited frequency content, such as 4 kHz audio in [90], extending the approach to more diverse music tasks requiring a wider frequency range, as well as to the bioacoustic domain requiring focus on specific frequency ranges depending on the task, remains a future research direction.
The CLAP features of M2D-CLAP also have a limitation in the data coverage for general-purpose usage. Achieving general-purpose performance, as in cutting-edge LLM, may require a similar scale of training data. To this end, further advancement of approaches such as automated generation of audio captions, as in Auto-ACD[46], is necessary to address the challenge of data scalability.